### v1

@@ 1 Introduction
The scale $r_{\rm qu}$ at which non-abelian effects kick is the scale at which the system enters the ensemble of black hole states. As we argue in section 2, the latter scale agrees with the quantum scale of BPS $AdS_{2}$ black holes Preskill:1991tb ; Lin:2022zxd ; Heydeman:2020hhw ; Lin:2022rzw  – the $AdS_{2}$ throat depth at which quantum gravity fluctuations cannot be ignored, and modify the classical picture of the geometry. In the process of collapse of the near-BPS brane star, the constituent branes shrink their extent in the transverse space of the fivebranes; the cap in the geometry descends the $AdS_{2}$ throat until it reaches the quantum scale $r_{\rm qu}$, at which point the little string excitations are liberated and the system enters the black hole phase.
@@
This picture of a black hole interior cannot be solely that of vacuum solutions of general relativity. The inner horizon, even in general relativity, is expected to be the locus of large tidal forces for an infalling probe Marolf:2011dj ; Horowitz:2022leb , and the point where the geometry caps off. For this reason, it is natural to suspect that in a complete theory, there is structure at the quantum scale of an extremal horizon, where inner and outer horizons coincide, which probes go “splat” upon. It is less clear whether the outer “horizon” (the surface of the ultracompact object) is a hard surface away from extremality, that probes go “splat” upon. Non-extremality separates the inner and outer horizon (to the extent a geometrical picture of the black hole interior exists). If the surface of the non-extremal, collapsed brane star is simply a crossover between confined and deconfined phases of the little string, a highly blueshifted infalling probe may simply make a “jet” of fractionated constituents as it enters the star. Due to that fractionation, the probe is completely trapped once it passes the “horizon” (or more generally passes inside a “trapped surface” which extends out beyond the star’s surface to meet it). It necessarily thermalizes into the non-abelian soup in the interior, perhaps suddenly upon reaching the inner horizon, to be re-emitted as Hawking radiation. The region between inner and outer “horizons”, if it exists, would simply be the “thermal atmosphere” of the little strings (see sections 2 and 6.2 of Martinec:2014gka ). In accordance with the fuzzball proposal, Hawking radiation would be the process of assembling singlet fundamental strings from their fractionated constituents in this atmosphere, and emitting them from the outer surface of the ultracompact object.
@@ 2.3 BPS black fivebranes and extremal BTZ: Scales and rescalings
Another route to this result begins with the effective 2d Jackiw–Teitelboim (JT) gravity theory that captures the gravitational dynamics of a nearly-$AdS_{2}$ black hole throat. As an $AdS_{2}$ black hole nears extremality, quantum fluctuations of JT gravity’s “Schwarzian modes” start to dominate. An analysis of the near-BPS dynamics Lin:2022rzw of JT gravity corroborates the estimate (2.43) of the length of the extremal throat in the quantum theory.
@@ 2.5 Scales in supertubes
We should also note a few other “quantum” scales that arise in the analysis of two-charge states. Suppose we partition the fivebrane excitation budget into the transverse excitations of the scalars $\mathsf{F}$} that described where the fivebrane sits in its transverse space, and the internal excitations of the fields ${\mathfrak{a}}^{\scriptscriptstyle AB}$ which are the BPS excitations of the internal gauge multiplet on the fivebranes, with total excitation levels $n_{p}^{\perp}$ and $n_{p}^{\parallel}$ respectively. The classical analysis of the two-charge states using the magnetic effective string action (2.10) is valid when the nonabelian excitations of the fivebrane are heavy and thus can be consistently decoupled. But how heavy are they?
@@
$\mathcal{S}_{\rm DBI}=\tau_{D1}\int\!e^{-\Phi}\sqrt{{\rm det}[G\!+\!(B\!-\!F)]}+\tau_{D1}\int\!C\wedge e^{B-F}~.$ (2.48)
@@
$M_{\rm D1}=\frac{\big|{\mathsf{F}}^{\flat}_{m}(v)-{\mathsf{F}}^{\flat}_{m^{\prime}}(v)\big|}{\ell_{na}^{2}}~.$ (2.50)
@@
$\tau^{~}_{\rm D2}=\frac{\big|\hat{\mathsf{F}}_{m}(v)-\hat{\mathsf{F}}_{m^{\prime}}(v)\big|}{4\pi^{2}(\alpha^{\prime})^{3/2}}~.$ (2.51)
@@
There is an important point to bear in mind in this discussion. The scale $\mu$ set by the fivebrane tension differs from the non-abelian scale at which non-abelian effects arise, by a factor $V_{4}/(\alpha^{\prime})^{2}]^{1/2}$. The non-abelian scale, where stretched D-branes become as light as fundamental strings is $|\delta\hat{\mathsf{F}}|^{2}=O(\alpha^{\prime})$, while the intrinsic width of the fivebrane center-of-mass wavefunction is $|\delta\hat{\mathsf{F}}|^{2}=O\big({\alpha^{\prime}}^{2}/V_{4}\big)$. Increasing $V_{4}$ makes the fivebranes heavier, and the size of the NS5-P supertube smaller. On the other hand, the energy cost of stretched D1- or D2-branes is independent of $V_{4}$. For $V_{4}$ larger than the string scale, non-abelian effects between a pair of fivebranes kick in somewhat before their center-of-mass wavefunctions overlap. In particular, the superstratum throat length (2.47) at which quantum effects take over is not where $J_{R}\sim O(1)$, but rather $J_{R}\sim O\big(V_{4}/{\alpha^{\prime}}^{2}\big)$. With this caveat, we will treat these two scales as roughly interchangeable, keeping in mind that if the torus is significantly larger than string scale, quantum effects kick in somewhat earlier than fivebrane separations of order $\mu$.
@@ 3.3 Layer 2
As discussed in the introduction, the singularities at ${\mathsf{x}}={\mathsf{F}}_{\mathsf{m}}$ are resolved by $\alpha^{\prime}$ effects – string wavefunctions don’t fit down isolated fivebrane throats. This phenomenon was originally observed in a particular solvable example of separated NS5-branes (in the absence of string or momentum charge), where the fivebranes are placed in a ${\mathbb{Z}}_{n}$-symmetric array on a circle of radius $\mathfrak{a}$} on their Coulomb branch. In this special source configuration, worldsheet string theory is exactly solvable in the decoupling limit (1.2Giveon:1999px ; Giveon:1999tq , and the loop expansion is well-behaved provided the radius $\mathfrak{a}$} is sufficiently large (as discussed in the introduction and in section 2.5, sufficiently large that D-branes stretched between the fivebranes are sufficiently heavy and thus suppressed). In the construction of Giveon:1999px ; Giveon:1999tq , the worldsheet conformal field theory is a gauged Wess-Zumino-Witten (WZW) model for the coset
@@ 3.5 Conserved charges
$g_{vi}=-\frac{\omega_{i}}{Z_{1}}\sim-\frac{2\kappa_{D}^{2}}{\Omega_{D-2}}\bigg(\frac{\mathcal{P}^{i}(v)}{r^{D-3}}-\frac{\mathcal{J}^{ij}{\mathsf{x}}^{j}}{r^{D-1}}\bigg)~,$ (3.69)
@@
$\qquad-\frac{1}{4\pi}\sum_{{\mathsf{m}}=1}^{{n_{5}}}\sum_{{\mathsf{n}}=1}^{{n_{1}}}\frac{{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},i}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},j}-{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},j}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},i}+\epsilon_{ijkl}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}}{|{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}|^{2}}~.$ (3.74)
@@ 5 The decoupling limit
As in section 2.3, will be interested in taking the fivebrane decoupling limit by making a similar rescaling of the transverse space coordinates (along the lines of (1.2)) [^14] We do not include factors of $n_{1}n_{5}$ or $R_{y}$ as in (2.22), which are specific to the $AdS_{3}$ decoupling limit considered there.
@@
$\hat{\ell}_{na}=\bigg[\frac{V_{4}}{(\alpha^{\prime})^{2}}\bigg]^{\frac{1}{2}}~.$ (5.3)
@@
$\qquad-\sum_{{\mathsf{m}}=1}^{{n_{5}}}\sum_{{\mathsf{n}}=1}^{{n_{1}}}\frac{{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{i}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{j}-{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{j}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{i}+\epsilon_{ijkl}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{k}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{l}}{|{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}|^{2}}\Bigg)~.$ (5.10)
@@ 6 Circular superstrata and their stringy siblings
When the charge quanta have the hierarchy $n_{p}<{n_{1}}$ in the original duality frame of P excitations on top of an NS5-F1 supertube, this T-dual frame (where the sources are NS5-P and F1) is in fact the appropriate one for describing the near-source structure of superstrata in this regime of charge quanta. It was pointed out in Martinec:2023xvf that the NS5-P frame describes the near-source structure of generic two-charge BPS geometries. For small enough values of the third (F1) charge in this duality frame, the geometry is not sufficiently deformed by the backreaction of the string winding to change this conclusion, and so continues to be the appropriate frame to describe the three-charge near-source structure. The geometry will be that of Eq. (2.33) until one reaches the outer envelope of the fivebrane star and beyond out to the radius $r_{d}=1/\sqrt{{n_{5}}}$ of Eq. (2.31).
@@ 6.1 Round source profiles
Evaluating the conserved charges $n_{p},J_{L},J_{R}$ associated to the currents (5.8), (5) of this background, one finds
$n_{p}$ $=\frac{{\mathsf{p}}^{2}{\mathfrak{a}}^{2}}{n_{5}\mu^{2}}+\frac{{n_{1}}}{{n_{5}}w_{y}^{2}|{\mathfrak{a}}^{2}-{\mathfrak{b}}^{2}|}\Big({\mathsf{p}}^{2}{\mathfrak{a}}^{2}w_{y}^{2}+{n_{5}}^{2}k^{2}{\mathfrak{b}}^{2}-{n_{5}}\,k\,{\mathsf{p}}w_{y}\times\text{min}({\mathfrak{a}}^{2},{\mathfrak{b}}^{2})\Big)~~,$
@@
$J_{L}$ $=\frac{{\mathsf{p}}{\mathfrak{a}}^{2}}{2\mu^{2}}-\frac{{n_{1}}}{w_{y}|{\mathfrak{a}}^{2}-{\mathfrak{b}}^{2}|}\Big({\mathsf{p}}w_{y}{\mathfrak{a}}^{2}+{n_{5}}k{\mathfrak{b}}^{2}-({n_{5}}\,k+{\mathsf{p}}w_{y})\times\text{min}({\mathfrak{a}}^{2},{\mathfrak{b}}^{2})\Big)~.$
@@ 6.2 Comparing to single-mode effective superstrata
We can compare the preceding solution to the effective description of round supestrata derived in Bena:2025uyg . These solutions have, in addition to the NS-NS fields we have restricted our consideration to, a profile ${\mathfrak{a}}_{0}(v)$ of the R-R scalar component of the tensor multiplet on the fivebrane, which sources R-R fields in the bulk (see Martinec:2024emf for a discussion in the present context). However, in the analysis Bena:2025uyg , the polarization state of the momentum excitation was largely irrelevant – only the overall momentum flux mattered when the solutions were coarse-grained by averaging over their oscillations along the $y$-circle. We will thus proceed, under the assumption that the similarities with our solutions largely outweigh the differences. As mentioned above, this seems a reasonable assumption, since in the WKB limit we employ, the polarization state of the F1-P sources in our solution is a minor effect compared to the overall momentum/winding carried by the strings.
@@
processed through smearing and the T-duality transformation (6.2). We will identify the parameter $a$ of superstrata with the parameter $\frac{\alpha^{\prime}}{R_{\tilde{y}}^{2}}{\mathfrak{a}}$ of our circular source solutions in Eq. (6.41) below.
@@
where we have dropped the subscript labels ${\mathsf{m}},{\mathsf{n}}$ on the source locations ${\mathsf{F}},{\mathsf{X}}$ since there is only one kind of source of each type. These relations has been corroborated via an analysis of worldsheet string theory in circular supertube background in the NS5-F1 duality frame Martinec:2025xoy ; Bena:2025uyg ; we repeat this analysis in the T-dual NS5-P frame in Appendix C. Here $\ell,m,n$ are the center-of-mass quantum numbers of each of the momentum quanta in the coherent condensate at radius $|{\mathsf{x}}_{P}|=\hat{\mathsf{b}}$ (which under T-duality become wound strings located at $|{\mathsf{X}}|={\mathfrak{b}}$).
@@
We thus find complete agreement for between the smeared momentum harmonic function (when the strings themselves carry no momentum), and the string winding harmonic function of the effective superstratum.

### v2

@@ 1 Introduction
The scale $r_{\rm qu}$ at which non-abelian effects kick in is the scale at which the system enters the ensemble of black hole states. As we argue in section 2, this non-abelian scale agrees with the quantum scale of BPS $AdS_{2}$ black holes Preskill:1991tb ; Lin:2022zxd ; Heydeman:2020hhw ; Lin:2022rzw  – the $AdS_{2}$ throat depth at which quantum gravity fluctuations cannot be ignored, and modify the classical picture of the geometry. In the process of collapse of the near-BPS brane star, the constituent branes shrink their extent in the transverse space of the fivebranes; the cap in the geometry descends the $AdS_{2}$ throat until it reaches the quantum scale $r_{\rm qu}$, at which point the little string excitations are liberated and the system enters the black hole phase.
@@
This picture of a black hole interior cannot be solely that of vacuum solutions of general relativity. The inner horizon, even in general relativity, is expected to be the locus of large tidal forces for an infalling probe Marolf:2011dj ; Horowitz:2022leb , and the point where the geometry caps off. For this reason, it is natural to suspect that in a complete theory, there is structure at the quantum scale of an extremal horizon, where inner and outer horizons coincide, which probes go “splat” upon. It is less clear whether the outer “horizon” (the surface of the ultracompact object) is a hard surface away from extremality, that probes go “splat” upon. Non-extremality separates the inner and outer horizon (to the extent a geometrical picture of the black hole interior exists). If the surface of the non-extremal, collapsed brane star is simply a crossover between confined and deconfined phases of the little string, then pursuing the bag model analogy, a highly blueshifted infalling probe may simply make a “jet” of fractionated constituents as it enters the star. Due to that fractionation, the probe is completely trapped once it passes the “horizon” (or more generally passes inside a “trapped surface” which extends out beyond the star’s surface to meet it). It necessarily thermalizes into the non-abelian soup in the interior, perhaps suddenly upon reaching the inner horizon, to be re-emitted as Hawking radiation. The region between inner and outer “horizons”, if it exists, would simply be the “thermal atmosphere” of the little strings (see sections 2 and 6.2 of Martinec:2014gka ). In accordance with the fuzzball proposal, Hawking radiation would be the process of assembling singlet fundamental strings from their fractionated constituents in this atmosphere, and emitting them from the outer surface of the ultracompact object.
@@ 2.3 BPS black fivebranes and extremal BTZ: Scales and rescalings
Another route to this result begins with the effective 2d Jackiw–Teitelboim (JT) gravity theory that captures the gravitational dynamics of a nearly-$AdS_{2}$ black hole throat. As an $AdS_{2}$ black hole nears extremality, quantum fluctuations of JT gravity’s “Schwarzian modes” start to dominate. An analysis of the near-BPS dynamics as well as the wavefunction of the BPS ground state of JT supergravity Lin:2022rzw corroborates the estimate (2.43) of the length of the extremal throat in the quantum theory.
@@ 2.5 Scales in supertubes
We should also note a few other quantum scales, that arise in the analysis of two-charge states. Suppose we partition the fivebrane excitation budget into the transverse excitations of the scalars $\mathsf{F}$} that described where the fivebrane sits in its transverse space, and the internal excitations of the fields ${\mathfrak{a}}^{\scriptscriptstyle AB}$ which are the BPS excitations of the internal gauge multiplet on the fivebranes, with total excitation levels $n_{p}^{\perp}$ and $n_{p}^{\parallel}$ respectively. The classical analysis of the two-charge states using the magnetic effective string action (2.10) is valid when the nonabelian excitations of the fivebrane are heavy and thus can be consistently decoupled. But how heavy are they?
@@
$\mathcal{S}_{\rm DBI}=\tau_{D1}\int\!e^{-\Phi}\sqrt{{\rm det}\big[G\!+\!(B\!-\!F)\big]}+\tau_{D1}\int\!C\wedge e^{B-F}~.$ (2.48)
@@
$M_{\rm D1}=\frac{\big|{\mathsf{F}}^{\flat}_{\mathsf{m}}(v)-{\mathsf{F}}^{\flat}_{{\mathsf{m}}^{\prime}}(v)\big|}{\ell_{na}^{2}}~.$ (2.50)
@@
$\tau^{~}_{\rm D2}=\frac{\big|\hat{\mathsf{F}}_{\mathsf{m}}(v)-\hat{\mathsf{F}}_{{\mathsf{m}}^{\prime}}(v)\big|}{4\pi^{2}(\alpha^{\prime})^{3/2}}~.$ (2.51)
@@
There is an important point to bear in mind in this discussion. The scale $\mu$ set by the fivebrane tension differs from the non-abelian scale at which non-abelian effects arise, by a factor $V_{4}/(\alpha^{\prime})^{2}]^{1/4}$. The non-abelian scale, where stretched D-branes become as light as fundamental strings is $|\delta\hat{\mathsf{F}}|^{2}=O(\alpha^{\prime})$, while the intrinsic width of the fivebrane center-of-mass wavefunction is $|\delta\hat{\mathsf{F}}|^{2}=O\big({\alpha^{\prime}}/(V_{4})^{\frac{1}{2}}\big)$. Increasing $V_{4}$ makes the fivebranes heavier, and the size of the NS5-P supertube smaller. On the other hand, the energy cost of stretched D1- or D2-branes is independent of $V_{4}$. For $V_{4}$ larger than the string scale, non-abelian effects between a pair of fivebranes kick in somewhat before their center-of-mass wavefunctions overlap. In particular, the superstratum throat length (2.47) at which quantum effects take over is not where $J_{R}\sim O(1)$, but rather $J_{R}\sim O\big((V_{4})^{\frac{1}{2}}/{\alpha^{\prime}}\big)$. With this caveat, we will treat these two scales as roughly interchangeable, keeping in mind that if the torus is significantly larger than string scale, quantum effects kick in somewhat earlier than fivebrane separations of order $\mu$.
@@ 3.3 Layer 2
As discussed in the introduction, the singularities at ${\mathsf{x}}={\mathsf{F}}_{\mathsf{m}}$ are resolved by $\alpha^{\prime}$ effects – string wavefunctions don’t fit down isolated fivebrane throats. This phenomenon was originally observed in a particular solvable example of separated NS5-branes (in the absence of string or momentum charge), where the fivebranes are placed in a ${\mathbb{Z}}_{n_{5}}$-symmetric array on a circle of radius $\mathfrak{a}$} on their Coulomb branch. In this special source configuration, worldsheet string theory is exactly solvable in the decoupling limit (1.2Giveon:1999px ; Giveon:1999tq , and the loop expansion is well-behaved provided the radius $\mathfrak{a}$} is sufficiently large (as discussed in the introduction and in section 2.5, sufficiently large that D-branes stretched between the fivebranes are sufficiently heavy and thus suppressed). In the construction of Giveon:1999px ; Giveon:1999tq , the worldsheet conformal field theory is a gauged Wess-Zumino-Witten (WZW) model for the coset
@@ 3.5 Conserved charges
$g_{vi}=-\frac{\omega_{i}}{Z_{1}}\sim-\frac{2\kappa_{D}^{2}}{\Omega_{D-2}}\bigg(\frac{\mathcal{P}^{i}(v)}{r^{D-3}}+\frac{\mathcal{J}^{ij}{\mathsf{x}}^{j}}{r^{D-1}}\bigg)~,$ (3.69)
@@
$\qquad+\frac{1}{4\pi}\sum_{{\mathsf{m}}=1}^{{n_{5}}}\sum_{{\mathsf{n}}=1}^{{n_{1}}}\frac{{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},i}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},j}-{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},j}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},i}+\epsilon_{ijkl}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},k}}{|{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}},l}|^{2}}~.$ (3.74)
@@ 5 The decoupling limit
As in section 2.3, we will be interested in taking the fivebrane decoupling limit by making a similar rescaling of the transverse space coordinates (along the lines of (1.2)) [^14] We do not include factors of $n_{1}n_{5}$ or $R_{y}$ as in (2.22), which are specific to the $AdS_{3}$ decoupling limit considered there.
@@
$\hat{\ell}_{na}^{2}=\bigg[\frac{V_{4}}{(\alpha^{\prime})^{2}}\bigg]^{\frac{1}{2}}~.$ (5.3)
@@
$\qquad+\sum_{{\mathsf{m}}=1}^{{n_{5}}}\sum_{{\mathsf{n}}=1}^{{n_{1}}}\frac{{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{i}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{j}-{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{j}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{i}+\epsilon_{ijkl}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{k}\partial_{v}{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}^{l}}{|{\mathsf{D}}_{{\mathsf{m}}{\mathsf{n}}}|^{2}}\Bigg)~.$ (5.10)
In terms of these, one has the angular momentum of left- and right-handed rotations of the angular three-sphere
$J_{L}=\mathcal{J}_{12}+\mathcal{J}_{34}~~,~~~~J_{R}=\mathcal{J}_{12}-\mathcal{J}_{34}~.$ (5.11)
@@ 6 Circular superstrata and their stringy siblings
When the charge quanta have the hierarchy $n_{p}<{n_{1}}$ in the original duality frame of P excitations on top of an NS5-F1 supertube, this T-dual frame (where the sources are NS5-P and F1) is in fact the appropriate one for describing the near-source structure of superstrata in this regime of charge quanta. It was pointed out in Martinec:2023xvf that the NS5-P frame describes the near-source structure of generic two-charge BPS geometries. For small enough values of the third (F1) charge in this duality frame, the geometry is not sufficiently deformed by the backreaction of the string winding to change this conclusion, and so continues to be the appropriate frame to describe the three-charge near-source structure. The geometry will be that of Eq. (2.33) until one reaches the outer envelope of the fivebrane star and beyond.
@@ 6.1 Round source profiles
Evaluating the conserved charges $n_{p},J_{L},J_{R}$ associated to the currents (5.8), (5) and (5.11) of this background, one finds
$n_{p}$ $=\frac{{\mathsf{p}}^{2}{\mathfrak{a}}^{2}}{n_{5}\mu^{2}}+\frac{{n_{1}}}{{n_{5}}w_{y}^{2}|{\mathfrak{b}}^{2}-{\mathfrak{a}}^{2}|}\Big({\mathsf{p}}^{2}{\mathfrak{a}}^{2}w_{y}^{2}+{n_{5}}^{2}k^{2}{\mathfrak{b}}^{2}-{2\,}{n_{5}}k\,{\mathsf{p}}w_{y}\times\text{min}({\mathfrak{a}}^{2},{\mathfrak{b}}^{2})\Big)~~,$
@@
$J_{L}$ $=\frac{{\mathsf{p}}{\mathfrak{a}}^{2}}{2\mu^{2}}+\frac{{n_{1}}}{w_{y}|{\mathfrak{b}}^{2}-{\mathfrak{a}}^{2}|}\Big({\mathsf{p}}w_{y}{\mathfrak{a}}^{2}+{n_{5}}k{\mathfrak{b}}^{2}-({n_{5}}\,k+{\mathsf{p}}w_{y})\times\text{min}({\mathfrak{a}}^{2},{\mathfrak{b}}^{2})\Big)~.$
@@ 6.2 Comparing to single-mode effective superstrata
We can compare the preceding solution to the effective description of round superstrata derived in Bena:2025uyg . These solutions have, in addition to the NS-NS fields we have restricted our consideration to, a profile ${\mathfrak{a}}_{0}(v)$ of the R-R scalar component of the tensor multiplet on the fivebrane, which sources R-R fields in the bulk (see Martinec:2024emf for a discussion in the present context). However, in the analysis Bena:2025uyg , the polarization state of the momentum excitation was largely irrelevant – only the overall momentum flux mattered when the solutions were coarse-grained by averaging over their oscillations along the $y$-circle. We will thus proceed, under the assumption that the similarities with our solutions largely outweigh the differences. As mentioned above, this seems a reasonable assumption, since in the WKB limit we employ, the polarization state of the F1-P sources in our solution is a minor effect compared to the overall momentum/winding carried by the strings.
@@
processed through smearing and the T-duality transformation (6.2). We will identify the parameter $a$ of superstrata with the parameter $x\frac{\alpha^{\prime}}{R_{\tilde{y}}^{2}}]^{\frac{1}{2}}{\mathfrak{a}}$ of our circular source solutions in Eq. (6.42) below; and of course the pitches match,
${\mathsf{p}}=\kappa~.$ (6.32)
@@
where we have dropped the subscript labels ${\mathsf{m}},{\mathsf{n}}$ on the source locations ${\mathsf{F}},{\mathsf{X}}$ since there is only one kind of source of each type (also, note that the momentum and angular momeentum quantum numbers $m,n$ are distinct from these omitted labels ${\mathsf{m}},{\mathsf{n}}$). These relations are consequences of the “bubble equations” of 5d supergravity Bena:2025uyg ; they can also be independently derived via an analysis of worldsheet string theory in circular supertube background in the NS5-F1 duality frame Martinec:2025xoy ; Bena:2025uyg ; we repeat this analysis in the T-dual NS5-P frame in Appendix C. Here $\ell,m,n$ are the center-of-mass quantum numbers of each of the momentum quanta in the coherent condensate at radius $|{\mathsf{x}}_{P}|=\hat{\mathsf{b}}$ (which under T-duality become wound strings located at $|{\mathsf{X}}|={\mathfrak{b}}$).
@@
We thus find complete agreement between the smeared momentum harmonic function (when the strings themselves carry no momentum), and the string winding harmonic function of the effective superstratum.